# Nonasymptotic Analysis of Accelerated Methods With Inexact Oracle Under Absolute Error Bound

### v1

@@ 1.2 Related works
d’Aspremont [11] analyzes the performance of iFGM to minimize Liptshitz smooth objective function over compact convex sets $\mathcal{Q}$, with the inexact gradient satisfying
|~$\left|\left\langle\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x}),\mathbf{y}-\mathbf{z}\right\rangle\right|\leq\delta,\quad\forall\mathbf{x},\mathbf{y},\mathbf{z}\in\mathcal{Q},$
and shows the sequence generated by iFGM satisfies
$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|}{A_{K}}+3\delta.$
This inexactness assumption can be derived from the absolute error assumption for the bounded feasible region with diameter $D$ as
@@
|~$\left|\left\langle\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x}),\mathbf{y}-\mathbf{z}\right\rangle\right|\leq\bar{b}\left\|\mathbf{y}-\mathbf{z}\right\|\leq\bar{b}\max_{\mathbf{y},\mathbf{z}\in\mathcal{Q}}\left\|\mathbf{y}-\mathbf{z}\right\|=\bar{b}D\triangleq\delta,$
which shows the dependence of the error term on the diameter of the feasible region.
The notion of first-order $(\delta,L)$-oracle is proposed in [16], which removes the requirements of boundedness of the feasible set and differentiability of the objective function with a generalized definition of convexity and Lipschitz smoothness. More specifically, [16] assumes that the oracle provides an inexact function and subgradient pair $(\tilde{f}_{\delta_{\mathbf{x}}}(\mathbf{x}),\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x}))$ that satisfies
$0\leq f(\mathbf{y})-(\tilde{f}_{\delta_{\mathbf{x}}}(\mathbf{x})+\left\langle\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x}),\mathbf{y}-\mathbf{x}\right\rangle\leq\frac{L}{2}\left\|\mathbf{x}-\mathbf{y}\right\|^{2}+\delta_{\mathbf{x}}.$
@@
With $(\delta,L)$-oracle, [16] shows the sequence generated by iFGM satisfies
@@
This result allows different inexactness levels for different iterations. However, to use the above results for an oracle that satisfies the absolute error assumption, the boundedness of the feasible region is needed [16], i.e., an oracles with absolute error assumption $\left|\tilde{f}(\mathbf{x})-f(\mathbf{x})\right|\leq\Delta_{1}$ and ~$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq\Delta_{2}$, $\tilde{f}_{\delta_{\mathbf{x}}}=\tilde{f}(\mathbf{x})-\Delta_{1}-\Delta_{2}D$, ~$\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x})=\tilde{\nabla}f(\mathbf{x})$, and $\delta_{\mathbf{x}}=2\Delta_{1}+2\Delta_{2}D$ (with $D$ being the feasible region diameter) is a $(\delta,L)-$oracle.
To overcome the boundedness drawback of the $(\delta,L)$-oracle, [49] proposes the inexact first-order oracle of degree $q$ as
$0\leq f(\mathbf{y})-(f(\mathbf{x})+\left\langle\tilde{\nabla}f_{\delta}(\mathbf{x}),\mathbf{y}-\mathbf{x}\right\rangle\leq\frac{L}{2}\left\|\mathbf{y}-\mathbf{x}\right\|^{2}+\delta\left\|\mathbf{x}-\mathbf{y}\right\|^{q},$
with $q\in[0,2)$. With the appropriate selection of the $q$ and $\delta$ parameters, any inexact gradient that satisfies the absolute error condition satisfies the right-hand-side inequality above without the requirement of the bounded feasible region. Under this new oracle assumption, the sequence generated by iFGM satisfies
$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{4(L+q\rho)\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{(K+1)(K+2)}+\frac{(K+3)(2-q)\delta^{\frac{2}{2-q}}}{2\rho^{\frac{q}{2-q}}},$
with $\rho>0$ be any fixed number. Minimizing the right-hand-side with respect to $\rho$, the convergence rate becomes
$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{4L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{(K+1)(K+2)}+\frac{8^{\frac{q}{2}}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{q}(K+3)}{((K+1)(K+2)(K+3))^{\frac{q}{2}}}\delta.$
This latter condition allows unbounded regions; however, its left-hand-side inequality requires the inexact gradient to be a subgradient, which can not be guaranteed under the absolute error assumption. Furthermore, in the convergence bound, the error term depends on the initial condition. Finally, the analysis is provided for fixed inexactness levels.
Another accelerated algorithm studied under inexact oracles is the Similar Triangles Method (STM) [27]– see the inexact version in Algorithm 2. Compared to FGM, it only requires one projection for each iteration and, hence, it is preferable for constrained problems. The convergence properties of iSTM are studied under both absolute and relative errors.
Algorithm 2 Inexact Similar Triangles Method (iSTM)
0:  $\mathbf{z}_{-1}=\mathbf{x}_{0}\in\mathbb{R}^{d}$, $\alpha_{-1}=A_{-1}=0$.
1:  for $k=0,\cdots,K$ do
2:     $A_{k}=A_{k}+\alpha_{k}$ with $\alpha_{k}^{2}\leq A_{k}$
3:     $\mathbf{x}_{k}=\frac{1}{A_{k}}(A_{k-1}\mathbf{y}_{k-1}+\alpha_{k}\mathbf{z}_{k-1})$
4:     $\mathbf{z}_{k}=\Pi_{\mathcal{Q}}(\mathbf{x}_{0}-\sum_{i=0}^{k}\frac{\alpha_{i}}{L}\tilde{\nabla}f(\mathbf{x}_{i}))$
5:     $\mathbf{y}_{k}=\frac{1}{A_{k}}(A_{k-1}\mathbf{y}_{k-1}+\alpha_{k}\mathbf{z}_{k})$
6:  end for
When the feasible region $\mathcal{Q}=\mathbb{R}^{d}$, i.e., no projection is required, iSTM can be formulated as iFGM. From $A_{k}=A_{k-1}+\alpha_{k}$, we can see that the $\mathbf{x}_{k}$ updates in iFGM and iSTM are the same with different indices since iFGM updates $\mathbf{x}_{k}$ at the end of each iteration while iSTM updates it at the beginning of each iteration. The $\mathbf{z}_{k}$ updates for the two algorithms are the same. Hence, we only need to consider the $\mathbf{y}_{k}$ updates. From step 3 of iSTM, $A_{k-1}\mathbf{y}_{k-1}=A_{k}\mathbf{x}_{k}-\alpha_{k}\mathbf{z}_{k-1}$. Inserting this identity into step 5, and by the update of step 4, we have
$\mathbf{y}_{k}=\mathbf{x}_{k}+\frac{\alpha_{k}}{A_{k}}(\mathbf{z}_{k}-\mathbf{z}_{k-1})=\mathbf{x}_{k}-\frac{\alpha_{k}^{2}}{A_{k}L}\tilde{\nabla}f(\mathbf{x}_{k}).$
When $\alpha_{k}^{2}=A_{k}$, the above update is equivalent to the step 2 in iFGM.
Furthermore, [2] analyzes iSTM under the inexact gradient with absolute error bound. The sequence generated by iSTM with $\alpha_{k}^{2}=A_{k},\forall k$, satisfies
$f(\mathbf{y}_{K})-f_{*}\leq\frac{8L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{K^{2}}+3\tilde{R}_{K}\delta+\frac{K\delta^{2}}{2L},$
with $\tilde{R}_{K}\triangleq\max_{0\leq j\leq K}\{\left\|\mathbf{z}_{j}-\mathbf{x}_{*}\right\|,\left\|\mathbf{x}_{j}-\mathbf{x}_{*}\right\|,\left\|\mathbf{y}_{j}-\mathbf{x}_{*}\right\|\}$. This convergence bound is only given under fixed inexactness levels and restricts the stepsize as $\alpha_{k}^{2}=A_{k}$. Another drawback is that $\tilde{R}_{K}$ is not exactly quantifiable; indeed, [2] only shows its boundedness by $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$ under the exact gradient oracle.
The boundedness of $\tilde{R}_{K}$ is shown by the same group under the relative error assumption with a relaxed stepsize [44]. It is shown that the sequence generated by iSTM with $\alpha_{k}=\frac{(k+2)^{p-1}}{2\alpha}$, $p\in[1,2]$, and $\alpha=\mathcal{O}(\max\{1,K^{\frac{p}{4}}\sqrt{\delta},K^{\frac{p}{2}}\delta,K^{p}\delta^{2}\})$ satisfies
$f(\mathbf{y}_{K})-f_{*}\leq\frac{16\alpha L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{(K+2)^{p}}.$
The paper shows $\tilde{R}_{K}$ is bounded by $2\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$ when the stepsize is properly chosen and, hence, there is no separate error term in the convergence rate.
Besides the abovementioned works, several researchers have proposed alternative error assumptions and results. The result of iFGM for the strongly convex setup is provided in [17]. [8] considers a variation of iFGM with the absolute error setting. To exploit the tradeoff between the convergence rate and accumulated error, the intermediate gradient method is proposed in [18] which is closely related to iFGM. This method is further generalized for other setups in [25, 19, 60, 28, 61].
[26] and [3] analyze nonaccelerated and accelerated algorithms under relative and absolute errors, respectively, with the latter focusing on the tradeoff between robustness to error and convergence rate–see also [32]. [24, 45] assume the error to satisfy the ~$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq B\left\|\mathbf{x}-\mathbf{x}_{*}\right\|$ condition, which can be derived from the relative error bound. [24] provides results for stochastic proximal inexact gradient methods with and without acceleration for strongly convex problems, while [45] studies generalized monotone variational inequality problems. Other notable contributions include the analysis of absolute and relative inaccuracy for proximal point methods in [57], the study of proximal gradient methods with absolute error in [59], the inexact proximal gradient method for weakly convex functions with absolute error in [37], and minimization of the accumulated error when the total allocation of error-control resources is fixed [15].
1.3 Contributions
This paper employs Performance Estimation Problem (PEP) techniques to primarily analyze the convergence rate of the Generalized Optimized Gradient Method with an inexact gradient oracle (iGOGM). Our key contributions are as follows:
  • We derive an analytical solution to the relaxed PEP optimization problem for iGOGM under variable oracle inexactness levels. The approach used to find this solution inspires our theoretical proof of the convergence. Notably, this proof can be understood without prior knowledge of PEP techniques.
  • The derived convergence bound consists of two components: the exact convergence rate and the accumulated error. Significantly, the error term is independent of the initial condition $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$, and is determined solely by the Lipschitz constant and the step size.
  • We provide a comprehensive analysis of the convergence bound from two perspectives: i) We investigate the tradeoff between the convergence rate and the accumulated error. ii) We establish optimal inexactness levels that preserve the method’s convergence rate.
  • We also provide the PEP-inspired convergence bound for the Generalized Fast Gradient Method with inexact gradient oracle (iGFGM).
We also provide the codes for our numerical experiments as well as the verification of the algebraic steps in our proofs. These codes are available at https://github.com/Yin-LIU/Inexact-Acceleration-PEP.
Notations

### v2

@@ 1.4 Related work
Bounded Inner Product Error (BIE)  d’Aspremont [8] focuses on smooth convex optimization over a compact convex set $\mathcal{Q}$, with the inexact gradient satisfying
[…]
The BIE condition can be derived from the absolute error assumption when the feasible set is bounded. Specifically, if the gradient error satisfies $\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq b$ for some constant $b$, then
@@
$\left|\left\langle\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x}),\mathbf{y}-\mathbf{z}\right\rangle\right|\leq b\left\|\mathbf{y}-\mathbf{z}\right\|\leq b\max_{\mathbf{y},\mathbf{z}\in\mathcal{Q}}\left\|\mathbf{y}-\mathbf{z}\right\|=b\cdot D\triangleq\delta,$
where $D=\max_{\mathbf{y},\mathbf{z}\in\mathcal{Q}}\left\|\mathbf{y}-\mathbf{z}\right\|$ represents the diameter of the feasible set. This derivation highlights that the error bound $\delta$ is directly proportional to the size of the feasible region.
Under the BIE assumption, iFGM has the convergence rate of
$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{A_{K}}+3\delta.$
Note that the convergence rate is determined by two components: the first term decreases with the accumulation parameter $A_{K}$, and the second term is a constant that scales with the inner product error bound $\delta$.
Inexact First-order $(\delta,L)$ Oracle (IFO)  Devolder et al. [12] introduce the $(\delta,L)$ first-order oracle condition for problems with unbounded feasible set or those with nonsmooth objective functions. The inexactness of the oracle pair $(\tilde{f}_{\delta_{\mathbf{x}}}(\mathbf{x}),\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x}))$ is quantified by
[…]
which is a relaxation of the first-order convexity and Lipschitz smoothness conditions.
@@
Under the IFO assumption, Devolder et al. [12] shows that the sequence generated by iFGM satisfies
@@
This result is similar to that of (BIE) condition, i.e., a term decreasing with iteration and an accumulated error term. However, (IFO) condition also allows varying inexactness levels along the iterations. While (IFO) condition allows an unbounded feasible set, obtaining it from the absolute error condition requires bounded feasible sets. Specifically, for an oracle with absolute errors $\left|\tilde{f}(\mathbf{x})-f(\mathbf{x})\right|\leq\Delta_{1}$ and $\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq\Delta_{2}$. By defining $\tilde{f}_{\delta_{\mathbf{x}}}\triangleq\tilde{f}(\mathbf{x})-\Delta_{1}-\Delta_{2}D$ and $\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x})\triangleq\tilde{\nabla}f(\mathbf{x})$, where $\delta_{\mathbf{x}}=2\Delta_{1}+2\Delta_{2}D$ (with $D$ representing the feasible region’s diameter), it can be shown the oracle with (AE) condition is a $(\delta,L)$-oracle [12].
Inexact First-order $(\delta,L)$ Oracle of Degree $q$ (IFO-q) To be able to address oracle inexactness in minimization of smooth nonconvex or nonsmooth convex objectives, while sacrificing the inexact gradients to be subgradients, Nabou et al. [41] propose the IFO-q oracle with an additional degree parameter $q\in[0,2)$ as